the eccentricity of ellipse (x-3)^2 + (y-4)^2= y^2/9 is


a) (3)^1/2/2


b) 1/3


C) 1/3(2)^1/2


d) 1/(3)^1/2

Grade:10

1 Answers

Sharandeep singh
11 years ago

taking under root 

{ (x-3)^2 + (y-4)^2 }^1/2 = (1/3)y

 

which is of the form 

{ (x-a)^2 + (y-b)^2 }^1/2 = e( ax + by +c)/ ( a^2 +b^2 ) ^1/2

on comparing e=1/3